If a-b=3 and b-c=5  then the value of a^2+b^2+c^2-ab-bc-ca=
 

Grade:12th pass

1 Answers

Arun
6 years ago
We can write the expression as
(1/2)*[(a-b)^2 + (b-c)^2 +(c-a)^2]
Now put values
a-b = 3
b-c = 5
c-a = 1
Hwnce the rquired answer =
= (1/2)*[3^2 + 5^2+ 1^2]
1/2* 35
= 35/2